# Robust CATE Estimation Using Novel Ensemble Methods

### v1

@@ 3.3 Estimators
  • Accurate GLM (AGLM): This method utilizes Lasso regression to fit a piece-wise constant function, by first encoding each variable into nested bins [20].
@@
  • Boosted Regression Trees: A sum of trees, where each tree is fitted on the residuals of the previous one [21].
@@
  • Random Forests (RF): An average of trees constructed with stochastic sampling of features and sample to induce variability in the trees, which acts as a form of regularization [22].
@@
  • Bayesian Additive Regression Trees (BART): Inspired by ensemble methods, with boosting in particular, BART also trains a sum-of-trees model, with the addition of a Bayesian regularization prior which controls the parameters of that model [23, 24].
A subset of the tested configurations are presented in the results section. Results for additional models are available in B.
Intermediate analysis of the performance of different estimators in different scenarios, revealed that no single estimator performs well across all scenarios (compared to others), with each single estimator under-performing in at least one of the scenario types. This motivated the search for an ensemble method that could provide a more robust performance than any of the methods comprising it. We used two ensemble approaches - Stacked X-Learner which uses stacking to ensemble difference prediction models within a Meta-Learner, and Consensus Based Averaging (CBA) which uses a combination of CATE estimators with high agreement.
@@
  • Outcome Ensemble: We used stacking for the first level ($\mu_{0}$, $\mu_{1}$) and second level ($\tau_{0}$, $\tau_{1}$) of nuisance functions. The estimators used for stacking were elasticnet, RF, boosting and AGLM, which were all cross-fitted using 3-fold CV and then used as input for linear regression for the outcome on all of the training set.
  • Consensus Based Averaging (CBA): This method prioritizes models with higher agreement. Initially, predictions from each model are collected. Kendall’s Tau correlation coefficients between each pair of models are calculated, and the mean correlation coefficient for each model is determined. Models are then sorted based on their mean correlation values. The method identifies a subset of models that are in higher agreement by detecting the ”knee” in the sorted mean correlation curve. Models up to this knee point are selected, and their predictions are averaged to produce the final ensemble prediction, enhancing robustness and accuracy. For further details on the method, see appendix A.
4 Simulation Study Results
The simulation study included generating multiple synthetic scenarios with varying characteristics and levels of complexity. For each DGP, training sets of multiple sizes were randomized and all of the CATE estimation methods described in 3 were fit on each of the training sets, and then used to predict CATE on 5000 held-out evaluation patients. Each combination of scenario and training-set size was used to generate 50 random training datasets. The results reported are based on the evaluation set.
@@ 4 Simulation Study Results
The study results demonstrate that there is no single model that is dominant across scenarios. Some models are in perfect conjunction with the DGP of some of the scenarios and therefore have an inherent advantage. For instance, the linear model-based T-learner shows dominance in larger linear scenarios, as can be seen in Figures 1 and 2.
@@
Figure 1 shows the advantage of ensemble methods (CBA and Stacked-X Learner) in achieving competitive accuracy, achieving best or near-best performance in most tested scenarios. Figure 2 confirms this for the RoD as well.
Figure 1: Relative Access Median sRMSE
@@
median sRMSE (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median sRMSE, compared to the best median sRMSE for that scenario.
Figure 2: Relative Access Median RoD
@@
median RoD (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median RoD, compared to the best median RoD for that scenario.
@@ 4.1 Subgroup Evaluation
Comparing the accuracy of aggregated CATE estimation in potential subgroups supports the above findings, with the ensemble methods having good accuracy in most scenarios. Notably, in this aggregated perspective, the performance of Causal Forest are also very robust and stable, possibly due to the ”tree-like” structure of SGs used for evaluation.
Due to the overlap in patients between the potential subgroups as they were defined in section 3.2.1, the estimation errors for subgroups are not independent of each other. Nevertheless, this perspective gives intuition for the expected usefulness of an estimator for subgroup identification and evaluation.
Figure 3: Relative Access Median sRMSE on Subgroups
median sRMSE on subgroups (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median sRMSE, compared to the best median sRMSE for that scenario.
Figure 4: Relative Access Median RoD on Subgroups
median RoD (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median RoD, compared to the best median RoD for that scenario.
@@ 4.2 Discussion
The results of this study highlight the robustness of ensemble methods in CATE estimation, when predicting effects at both the individual patient level and in potential subgroup evaluations. We propose ensembles of two different approaches—a meta-learner with model-stacking as a base-learner, and a Consensus-Based Average of estimators, which identifies a subset of estimators with the highest agreement. Both ensemble approaches demonstrate superior performance over single models, across a wide range of scenarios.
Ensemble methods consistently achieve high accuracy and exhibit robust performance across diverse scenarios. They maintain competitive performance relative to the best individual models, even when the underlying data generating process is unknown.
Given the inherent variability in model performance and the unknown nature of the true DGP in practical settings, the robustness of ensemble methods is particularly valuable. The proposed ensembles provide reliable and accurate CATE estimations, maintaining consistent performance across diverse scenarios. This robustness is crucial for practical applications in clinical trials and personalized medicine, where accurate and stable treatment effect predictions are essential for effective decision-making.
Additional work should expand the discussion to other types of endpoints, namely binary and time-to-event endpoints, which are common in clinical research. Other ensembling approaches have been proposed in the literature for CATE estimation [26, 27, 29, 28], however their performance was only demonstrated on much larger datasets than the ones considered here. Future work will benchmark selected methods on a unified set of scenarios.
@@ Appendix A Consensus Based Average
Let $\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{M}]$ be the matrix of predictions, where $\mathbf{p}_{i}$ represents the predictions from the $i$-th model, and $M$ is the number of models. Kendall’s Tau correlation coefficient between models $i$ and $j$ is given by:
$\tau_{ij}=\mathrm{Kendall^{\prime}s\;Tau}(p_{i},p_{j}).$
@@
$\bar{\tau}_{i}=\frac{1}{M-1}\sum_{j\neq i}\tau_{ij}.$
Next, the models are sorted in descending order based on their mean correlation values. To identify a subset of models that are in higher agreement, the method detects the ”knee” in the sorted mean correlation curve, which is the point where the difference between consecutive mean correlation values is minimized. Let $\bar{\tau}$ be the sorted mean correlation values:
$\bar{\tau}=[\bar{\tau}_{(1)},\bar{\tau}_{(2)},...\bar{\tau}_{(M)}].$
The ”knee” is detected by finding the index $k$ where the difference between consecutive mean correlation values is minimized:
@@
$\mathcal{S}=\{p_{(1)},p_{(2)},\ldots,p_{(k)}\}.$
@@
$CBA=\frac{1}{k}\sum_{p_{i}\in S}p_{i}.$
@@ Appendix B Supplementary Results
Table 2: median RoD (over 50 simluated datasets)
Appendix C PD-L1-biological data generation model
  • Tumor Growth: The tumor growth is driven upward by the mutation burden, and downward by the post-treatment levels of CD8+ effector.
[…]

### v3

@@ 7.1 Methods
  • Accurate GLM (AGLM): This method utilizes Lasso regression to fit a piece-wise constant function, by first encoding each variable into nested bins (Fujita et al., 2020).
@@
  • Boosted Regression Trees: A sum of trees, where each tree is fitted on the residuals of the previous one (Chen and Guestrin, 2016).
@@
  • Random Forests (RF): An average of trees constructed with stochastic sampling of features and sample to induce variability in the trees, which acts as a form of regularization (Breiman, 2001).
@@
  • Bayesian Additive Regression Trees (BART): Inspired by ensemble methods, with boosting in particular, BART also trains a sum-of-trees model, with the addition of a Bayesian regularization prior which controls the parameters of that model (Chipman et al., 2012; Dorie et al., 2022).
All meta-learners and ensemble methods used in the simulation use logistic regression for the estimation of the propensity score. The meta-learners and base-learners used in the simulation are: X-RF, X-BART, X-AGLM, X-Boosting, S-BART, T-Linear, and DR-RF. The mentioned base-learners were used for the estimation of the CATE and the relevant nuisance parameters. Furthermore, we compare two versions of debiased CF, with and without honesty named H-CF and CF respectively.
The ensemble methods CBA, R-Stacking, T-Stacking and Causal-Stacking utilize X-RF, X-Boosting, X-AGLM, T-Elastic-net, H-CF and CF as the base CATE estimators. We avoid a rigid train/validation split, and instead opted for cross-fitting the base CATE estimators, and fitting the ensemble on the entirety of the data. R-Stacking, T-Stacking and Causal-Stacking employ RF for estimation of the nuisance functions $m,\mu_{0},\mu_{1}$ (with limited cross-validation for selection of the hyper-parameters). R-Stacking utilizes logistic-regression for the estimation of the propensity score.
@@
The estimators used for the stacked-X were Elastic-net, RF, boosting, and AGLM, which were all cross-fitted using 3-fold CV and then used as input for linear regression for the outcome on all of the training set. Results for additional models are available in appendix B.
7.2 Data Generation
The DGPs used in the simulation study are ACIC based, the linear, slightly non-linear (NL), highly NL and the PD-L1. The DGPs are described in section 5.
All DGPs were used to generate training sets of sizes that are representative of clinical trial datasets, ranging from 100 patients to 1000 patients. The performance was then evaluated on held-out test sets of 5000 patients. The only exception to this rule are the ACIC datasets which, due to their finite sample size, were tested on the remaining part of the dataset not used for training (between 3800 and 4550 patients). Each combination of scenario and training-set size was used to generate 50 random training datasets. The results reported are based on the evaluation set.
7.3 Results
@@
The study results demonstrate that there is no single model that is dominant across scenarios. Some models are in perfect conjunction with the DGP of some of the scenarios and therefore have an inherent advantage. For instance, the linear model-based T-Learner shows dominance in larger linear scenarios, as can be seen in Figures 2 and 3.
@@
In general, the ensemble method appear to perform better than any single learner. Figure 2 shows the advantage of ensemble methods (CBA and Stacked-X Learner) in achieving competitive accuracy, achieving best or near-best performance in most tested scenarios. Furthermore, these ensemble methods also outperform the other ensemble approaches, R-Stacking, T-Stacking and Causal-Stacking in most scenarios. Figure 3 confirms this for the RoD as well.
Figure 2: Relative Excess Median sRMSE
@@
median sRMSE (over 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative excess error, compared to that of the best model.
Figure 3: Relative Excess Median RoD
@@
median RoD (over 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative excess error, compared to that of the best model.
@@ 7.4 Subgroup Evaluation
Comparing the accuracy of aggregated CATE estimation in potential subgroups supports the above findings, with the ensemble methods having good accuracy in most scenarios. Stacked-X Learner generally outperforms other ensembles.
Due to the overlap in patients between the potential subgroups as they were defined in section 6.1, the estimation errors for subgroups are not independent of each other. Nevertheless, this perspective gives intuition for the expected usefulness of an estimator for subgroup identification and evaluation.
Figure 4: Relative Excess Median sRMSE on Subgroups
median sRMSE on subgroups (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative excess median sRMSE, compared to the best median sRMSE for that scenario.
Figure 5: Relative Excess Median RoD on Subgroups
median RoD (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative excess median RoD, compared to the best median RoD for that scenario.
@@ 7.5 Discussion
The results of this study highlight the robustness of ensemble methods in CATE estimation, when predicting effects at both the individual patient level and in potential subgroup evaluations. We propose ensembles of two different approaches — an X-Learner with model-stacking as a base-learner, and a Consensus-Based Average of estimators, which identifies a subset of estimators with the highest agreement. Both ensembles approaches demonstrate superior performance over single models and ensemble models, across a wide range of scenarios.
The stacked-X performed particularly well. We hypothesize that the performance is driven by trying to minimize losses that are observed. The rest of the CATE ensembles (T-Stacking, R-Stacking and Causal-Stacking), try to minimize losses that are dependent on some CATE estimation. Therefore, when the CATE estimation of validation set is poor, the resulting ensemble under-performs. Still, most ensemble methods studied consistently achieve high accuracy and exhibit robust performance across diverse scenarios compared to the single-model.
Given the inherent variability in model performance and the unknown nature of the true DGP in practical settings, the robustness of ensemble methods is particularly valuable. This work evaluates the gain achieved by ensemble methods in setting that are relevant to clinical trial analysis - both in terms of the DGPs used and the small sample sizes used to train the models. It confirms that ensembling improves robustness also in relatively small sample sizes, despite the inherent challenge of splitting the already small dataset. This robustness is crucial for practical applications in clinical trials and personalized medicine, where accurate and stable treatment effect predictions are essential for effective decision-making.
Additional work should expand the discussion to other types of endpoints, namely binary and time-to-event endpoints, which are common in clinical research. Another question of interest is whether the advantage of Stacked X-Learner over other ensembles is preserved in other scenarios, specifically in larger data-sets.
@@ Appendix A Consensus Based Average
Let $\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{K}]$ be the matrix of predictions, where $\mathbf{p}_{i}$ represents the predictions from the $i$-th model, and $M$ is the number of models. Kendall’s Tau correlation coefficient between models $i$ and $j$ is given by:
$\mathcal{K}_{ij}=\mathrm{Kendall^{\prime}s\;Tau}(\mathbf{p}_{i},\mathbf{p}_{j}).$
@@
$\bar{\mathcal{K}}_{i}=\frac{1}{K-1}\sum_{j\neq i}\mathcal{K}_{ij}.$
Next, the models are sorted in descending order based on their mean correlation values. To identify a subset of models that are in higher agreement, the method detects the ”knee” in the sorted mean correlation curve, which is the point where the difference between consecutive mean correlation values is minimized. Let $\bar{\mathcal{K}}$ be the sorted mean correlation values:
$\bar{\mathcal{K}}=[\bar{\mathcal{K}}_{(1)},\bar{\mathcal{K}}_{(2)},...\bar{\mathcal{K}}_{(M)}].$
The ”knee” is detected by finding the index $m$ where the difference between consecutive mean correlation values is minimized:
@@
$\mathcal{S}=\{\mathbf{p}_{(1)},\mathbf{p}_{(2)},\ldots,\mathbf{p}_{(m)}\}.$
@@
$\mathbf{p}_{\text{ensemble}}=\frac{1}{m}\sum_{\mathbf{p}_{i}\in S}\mathbf{p}_{i}.$
@@ Appendix B Supplementary Results
Table 2: median RoD (over 50 simulated datasets)